From mar@roe.ac.uk Mon Jun 22 13:04:20 2015 Date: Mon, 15 Jun 2015 12:21:51 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , dmr@roe.ac.uk, Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 5 June 2015 WFAU VDFS Science Archives progress meeting minutes, 5 June 2015 Present: RPB, MSH, MAR, NJC, RSC Apologies: STV, KTN, AL, ETWS, ACD, RGM DoNM: 10am Friday 19 June 2015 in the Villas meeting room, ACD in the chair (room is booked). Actions discharged this time: ----------------------------- ACTION: MAR to compile a new list of non-surveys for release. Discharged. Actions carried forward from 24/4/2015 meeting: ----------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this action still has low priority. ACTION: RSC/ETWS to see if we can get another mxODBC trial period set-up as the current one had expired (was RSC to test the new version of mxODBC) Continues. ACTION: NJC noted that VISTA mosaicing crashes with normal filtering, which he should fix. Continues. NJC had implemented a fix and checked it into SVN. However, when it was run 'in anger' it still crashed. NJC to investigate. ACTION: RPB to compile a Web page listing all the completed releases. Continues. RPB had made a start compiling this list. However, the exercise revealed that a number of databases had been misplaced and he was investigating their whereabouts. ACTION: NJC to produce a roadmap for future VSA releases in order to plan for hardware requirements. Continues, NJC had contacted the PIs and there had been some discussion at the recent VVV meeting. ACTION: NJC noted that he had received a report of a problem with the ESO VVV DR1 release which he should investigate. In one pointing the merged band catalogue contained fewer sources than expected. Continues. This action was low on NJC's priority list. Actions resulting from the VVV DR3 brainstorming meeting held on 24/2/15: ------------------------------------------------------------------------ ACTION: NJC to speak to the VVV PIs, notably Dante, about the schedule for and expected number of remaining VVV releases. Continues. NJC had spoken to the PIs but the results had been inconclusive; he should pursue the matter further. (Most of the discussion at the VVV meeting had been about VVVEXT.) ACTION: RPB to experiment with ingest via partitioned tables, along the lines of the SSIS paper. Continues. ACTION: RGM (was RSC, RGM): (i) RSC to write a succinct summary of the ingest and outgest problem that is the principal bottleneck with the VVV processing (briefly the ingest and outgest speed is limited to approx. 100 Mbit/sec for no apparent reason) and (ii) RGM to compose an e-mail message around this summary and send it to Ani Thakar. Continues. RSC had composed a succinct summary and sent it to RGM. RGM had emailed Ani Thakar but no reply to date. RGM would send a reminder to Ani. Still awaiting a response. ACTION: NJC to remove dependencies on old libraries. Continues. This was mainly done but NJC to check on what if anything was still needed for VOTable handling. Specific points and new actions: -------------------------------- Project management: * Nothing for the minutes. WFCAM, VISTA and VST updates: * Nothing for the minutes. Networking: * MSH noted there was a planned outage of UK light on Saturday 27th June. WSA/VSA/OSA/GES Operations: * Processed VISTA data for Feb/March 2015 from CASU would be slightly delayed. * VVV CU16 had finished. ACTION: RPB to backup VSAVVV and copy to ramses16 Hardware and Systems: * MSH reported a dead disk on ranses9 had been replaced. Software: * NJC reported work on the setup of deep products for UKIDSS and schemas for VVV PSF and difference image products. * RSC noted that the changes implemented for CU19 speed ups had all been committed. He'd also been looking into a parallel version of CU19. Review of Deadlines: * See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Discussion was left for a future meeting. Survey Data Release: * Non since the last meeting. Non-survey Data Release: * Releases were being attempted for 14A projects, some had crashed and NJC/RPB were investigating. Astrogrid deployment, VO & Data Analysis services: * MSH noted he'd generated GLIMPSE coverage maps. Miscellaneous: * RSC would sit-in on the ROE MOONS meeting.